# CASE: Efficient Curricular Data Pre-training for Building Assistive Psychology Expert Models

### v2

@@ Abstract
The limited availability of psychologists necessitates efficient identification of individuals requiring urgent mental healthcare. This study explores the use of Natural Language Processing (NLP) pipelines to analyze text data from online mental health forums used for consultations. By analyzing forum posts, these pipelines can flag users who may require immediate professional attention. A crucial challenge in this domain is data privacy and scarcity. To address this, we propose utilizing readily available curricular texts used in institutes specializing in mental health for pre-training the NLP pipelines. This helps us mimic the training process of a psychologist. Our work presents CASE-BERT that flags potential mental health disorders based on forum text. CASE-BERT demonstrates superior performance compared to existing methods, achieving an f1 score of 0.91 for Depression and 0.88 for Anxiety, two of the most commonly reported mental health disorders. Our code is publicly available  [^1] https://github.com/sarthakharne/CASE
@@
Sarthak Harne1, Monjoy Narayan Choudhury1, Madhav Rao1, TK Srikanth1, Seema Mehrotra2, Apoorva Vashisht2, Aarushi Basu3, Manjit Sodhi 4, 1IIIT Bangalore, 2NIMHANS, 3Ashoka University, 4IBM, Correspondence: sarthak.harne@iiitb.ac.in, monjoy.choudhury@iiitb.ac.in
@@
Sarthak Harne1, Monjoy Narayan Choudhury1, Madhav Rao1, TK Srikanth1, Seema Mehrotra2, Apoorva Vashisht2, Aarushi Basu3, Manjit Sodhi 4, 1IIIT Bangalore, 2NIMHANS, 3Ashoka University, 4IBM, Correspondence: sarthak.harne@iiitb.ac.in, monjoy.choudhury@iiitb.ac.in
@@ 1 Introduction
Our work builds on the philosophy of how a person gains curricular knowledge. Our approach is analogous to the training of a psychology student as mandated by the American Psychology Association (APA) APA (2019) from a theoretical standpoint. Building on previous work of Gunasekar et al. (2023) where they use high-quality data for building a small, yet effective generative model for Python Programming, we pre-train off-the-shelf models using high-quality curricular data provided by psychologists in academia. This high-quality data helps us to tackle domain data scarcity by using a small amount of data to pre-train and finetune state-of-the-art (SOTA) discriminative assistive psychology expert models. We believe our contribution in the form of curriculum learning for discriminative models can be used in other domains also, given the amount of curriculum data present in the domain is respectable.
@@ 2.2 Textbooks are all you need
Gunasekar et al. (2023) in their work introduce the notion of how the quality of data affects the performance of text generation of Large Language Models. In this work, they rely highly on the hypothesis that involves the usage of "textbook quality" data from the web (6B tokens) as well as synthetically generated data from GPT3.5. This small-scale data compared to large web-crawled corpora gave a respectable performance, especially for a significantly lower number of parameters. They pre-train on such data and then attempt to fine-tune on "textbook-exercise" like data. In this way they attempt to capture the domain learning process for humans. We aim to verify this ideology in the case of mental health disorder identification in discriminative models.
@@ 4.3 Prompting
Due to the recent availability of capable instruction-tuned LLMs, OpenAI (2023), MetaAI (2024) it has become easier to solve general problems via merely prompting these models Nazir and Wang (2023). To compare how well these discriminative models perform against modern generative LLMs, we test these LLMs on the same datasets. For this, we use a few-shot prompting strategy. The prompt structure used for inference on CounselChat Bertagnolli (2020) is shown in Fig.3. We pick a random instance from the train split of the dataset to be passed on as the one-shot example. Subsequently, we pass one instance from the test set for the model to infer.

### v3

@@ Abstract
The limited availability of psychologists necessitates efficient identification of individuals requiring urgent mental healthcare. This study explores the use of Natural Language Processing (NLP) pipelines to analyze text data from online mental health forums used for consultations. By analyzing forum posts, these pipelines can flag users who may require immediate professional attention. A crucial challenge in this domain is data privacy and scarcity. To address this, we propose utilizing readily available curricular texts used in institutes specializing in mental health for pre-training the NLP pipelines. This helps us mimic the training process of a psychologist. Our work presents CASE-BERT that flags potential mental health disorders based on forum text. CASE-BERT demonstrates superior performance compared to existing methods, achieving an f1 score of 0.91 for Depression and 0.88 for Anxiety, two of the most commonly reported mental health disorders. Our code and data are publicly available.  [^1] https://github.com/sarthakharne/CASE
@@
Sarthak Harne*1, Monjoy Narayan Choudhury*1, Madhav Rao1, TK Srikanth1, Seema Mehrotra2, Apoorva Vashisht2, Aarushi Basu3, Manjit Sodhi 4 1IIIT Bangalore, 2NIMHANS, 3Ashoka University, 4IBM, Correspondence: sarthak.harne@iiitb.ac.in, monjoy.choudhury@iiitb.ac.in *Equal Contribution.
@@
Sarthak Harne*1, Monjoy Narayan Choudhury*1, Madhav Rao1, TK Srikanth1, Seema Mehrotra2, Apoorva Vashisht2, Aarushi Basu3, Manjit Sodhi 4 1IIIT Bangalore, 2NIMHANS, 3Ashoka University, 4IBM, Correspondence: sarthak.harne@iiitb.ac.in, monjoy.choudhury@iiitb.ac.in *Equal Contribution.
@@ 1 Introduction
Our work builds on the philosophy of how a person gains curricular knowledge. Our approach is analogous to the training of a psychology student as mandated by the American Psychology Association (APA) APA (2019) from a theoretical standpoint. Building on Textbooks Are All You Need Gunasekar et al. (2023) where they use high-quality data for building a small, yet effective generative model for Python Programming, we pre-train off-the-shelf models using high-quality curricular data provided by psychologists in academia. This high-quality data helps us to tackle domain data scarcity by using a small amount of data to pre-train and finetune state-of-the-art (SOTA) discriminative assistive psychology expert models. We believe our contribution in the form of curriculum learning for discriminative models can be used in other domains also, given the amount of curriculum data present in the domain is respectable.
@@ 2.2 Textbooks Are All You Need
Textbooks Are All You Need Gunasekar et al. (2023) introduces the notion of how the quality of data affects the performance of text generation of Large Language Models. In this work, they rely highly on the hypothesis that involves the usage of "textbook quality" data from the web (6B tokens) as well as synthetically generated data from GPT3.5. This small-scale data compared to large web-crawled corpora gave a respectable performance, especially for a significantly lower number of parameters. They pre-train on such data and then attempt to fine-tune on "textbook-exercise" like data. In this way they attempt to capture the domain learning process for humans. We aim to verify this ideology in the case of mental health disorder identification in discriminative models.
@@ 4.3 Prompting
Due to the recent availability of capable instruction-tuned LLMs, OpenAI (2023), MetaAI (2024) it has become easier to solve general problems via merely prompting these models Nazir and Wang (2023). To compare how well these discriminative models perform against modern generative LLMs, we test these LLMs on the same datasets. For this, we use a few-shot prompting strategy. The prompt structure used for inference on CounselChat Bertagnolli (2020) is shown in Fig.3. We first experimented with various prompt styles and picked this prompt based on some subjective analysis. We choose a random instance from the train split of the dataset to be passed on as the one-shot example. Subsequently, we pass one instance from the test set for the model to infer.